I have a number with unit "s2/rad2/kg" and I want to convert it into "N/m", and I'm wondering how to do this...

The number is 0.06587 s2/rad2/kg
Answer #1

Solution: We know that Newton (N) is unit of force .The formula of force is given as follows :

Force = mass x acceleration

Where :

Unit of force =Newton (N)

Unit of mass = kilogram(kg )

Unit of acceleration= m/s2

Thus from above discussion it is clear that :

    Newton = kg - m /s2

=> N= kg - m /s2

  =>    N/m = kg/s2

=> s2/kg = m/ N ( since taking inverse both side ). =>  s2/ rad2/ kg = m/ rad2/ N ( since dividing both side by rad2)

Hence given number .06587 s2/rad 2/kg can be written , in required unit ,as .06587 m /rad 2/ N .